OpenStudy (anonymous):

Select the equation of the line in slope-intercept form: 4x + 9y = –36 (Points : 4) 9y = 4x + 36 9y = –4x – 36 y= -4/9 x - 4 y= 4/9x +4

OpenStudy (anonymous):

The graph of a line has a slope of 2 and passes through the point (–6, –5). What is the equation of the line in point-slope form? (Points : 4) y – 6 = 2(x – 5) y + 6 = 2(x + 5) y – 5 = 2(x – 6) y + 5 = 2(x + 6)

OpenStudy (anonymous):

@OpiGeode

OpenStudy (anonymous):

@Miss.SweetiePie